# bibliography.bib

### v1

@@
Abstract: Using cost-optimisation to support energy system design decisions hides from view many more feasible designs that, while economically comparable to the least-cost option, may have other practical advantages that stakeholders prefer. Increasingly, modellers look at Modelling to generate alternatives (MGA) as a way to go beyond few, optimal designs and provide stakeholders with broad portfolios of design options to appraise based on their preferences and knowledge of the problem. In practice, however, generating and exploring all the feasible system design options is not possible for finite computational power. Modellers need to choose what design feature to generate diversity around, and yet, they cannot know what features may matter the most to heterogeneous real-world stakeholders. Stakeholders’ knowledge is thus needed to guide the MGA search towards the most important trade-offs, but how to incorporate such knowledge in the MGA workflow is unclear. To address this gap, we propose a first-of-its-kind human-in-the-loop (HITL) approach that seamlessly and automatically integrates stakeholder preferences into an MGA workflow. In a nutshell, we elicit stakeholders’ system design preferences via a first interaction between them and an initial, tentative MGA design space. Hence, we decode those preferences so that it becomes possible to feed them back to the MGA algorithm to perform a guided search. This search produces an updated, human-trained design space with a higher share of designs that align with the elicited preferences. In a controlled experiment for the case study of the Portuguese energy system, we demonstrate that our HITL-MGA approach may facilitate the identification of a consensus solution that balances many conflicting stakeholder preferences. The approach shows promise for generalisation and real-world application to accelerate technically and socially feasible energy transition decisions.
@@ 1 Introduction
Despite the urgency of climate mitigation, progress on the energy transition remains insufficient [iea_breakthrough_2023]. One of the key reasons for such slow decision-making is the difficulty in finding carbon-neutral energy system options that are both technically feasible and acceptable for the many involved stakeholders, from citizens to industry, system operators, and policymakers [susser_why_2022]. Energy system optimisation models are increasingly looked at as a way to support these complex planning decisions, but severe limitations hamper their effectiveness. In particular, previous work has shown that the conventional practice of optimising the system design for cost hides from view many other feasible designs that are economically comparable – i.e., they are marginally more costly, or within the cost uncertainty range – while potentially having other un-modelled benefits that make them substantially more practically viable [lombardi_policy_2020].
@@
Approaches known as ‘Modelling to Generate Alternatives’, or MGA, first developed in the late 70s [brill_use_1979] make it possible to go beyond few, optimal solutions and enlarge the decision space to hundreds of different possibilities, giving room for stakeholders to discuss and identify a consensus solution [decarolis_using_2011, decarolis_modelling_2016]. They have thus seen a spike in popularity: in the last few years, novel versions of MGA have been proposed for the specific needs of state-of-the-art large-scale energy system optimisation models that, representing large geographical regions at a high spatial resolution, entail countless degrees of spatial and technological freedom to achieve a feasible system design. Such novel MGA versions aim, for instance, at generating distinct spatial configurations of infrastructure deployment [lombardi_policy_2020] or at providing more homogeneous samples of the very large space of feasible decisions [pedersen_modeling_2021, neumann_broad_2023]. Increasingly, MGA is recognised as a critical tool to deliver meaningful, more transparent results [pickering_diversity_2022]; however, there are computational trade-offs in the use of MGA. For models of very large size, it is not yet computationally affordable to generate all the relevant design alternatives, and MGA algorithms need to decide what kind of diversity to prioritise in the design space. For instance, they must decide whether to focus more on discovering alternative technology mixes or different spatial configurations of technology deployment, for a given available amount of computational power [lombardi_what_2023]. And yet, determining which design options are more worth focussing on, insofar as they may be more relevant than others in facilitating real-world consensus formation, is not trivial.
From this perspective, stakeholder engagement has untapped potential as a way to ensure that the customisation of the MGA algorithm’s search strategy goes in the direction required by real-world information needs [lombardi_what_2023]. A stakeholder-informed MGA search could enable a targeted, efficient use of limited computational power while simultaneously answering the increasing calls for tighter integration between scientific and practical knowledge in the co-creation of model results [mcgookin_advancing_2024]. However, what is yet to be understood is how exactly modellers may effectively incorporate stakeholders’ knowledge of the problem into an MGA-based modelling workflow.
@@
To test the above hypothesis, we perform an experiment under controlled conditions in a case study of the Portuguese energy system. For our experiment, we rely on the SPORES MGA algorithm [lombardi_policy_2020], which lends itself to the proposed guided search due to its high degree of customisability, as we highlighted in previous work [lombardi_what_2023]. The controlled conditions, represented by our use of synthetic stakeholder preferences, allow us to have a clear benchmark to quantitatively assess the relative merits of the updated, human-trained design space in terms of its degree of alignment with the assumed preferences and its capacity to facilitate consensus formation. Our results demonstrate that the novel HITL-MGA approach comes with the expected advantages, and shows promise for generalisability and real-world application as a means to support the acceleration of technically and socially feasible energy planning decisions.
@@ 2 Methods
2.1 HITL-MGA workflow and experimental setup
@@
Figure 1: Conceptual representation of the proposed HITL-MGA workflow and of the controlled experiment we carry out in this study to test the workflow. The theoretical framework we propose envisions that we collect the high-level system design preferences arising from a first iteration of an MGA design space with stakeholders. Then, we decode which technical features underlying such preferences and translate them into parameters that may inform and re-tune the MGA search for design options. The resulting updated design space is richer in design options that align with stakeholders’ high-level preferences, and it facilitates consensus formation. In our controlled experiment, we replace stakeholder interaction with synthetic high-level stakeholder preferences, from which we derive five top-rated designs to initiate the HITL-MGA workflow. This way, we can test the outputs of the HITL-MGA workflow against the initial high-level preferences in a way that would not be possible if interacting with real-world stakeholders whose high-level preferences are typically unknown.
@@
In this study’s controlled experiment, we test the method based on synthetic high-level stakeholder preferences, which allows us to benchmark the outputs in a way that would not be possible if interacting with real-world stakeholders whose high-level preferences, as anticipated above, may be unknown. We assume five high-level preferences:
@@
    A slow rate of deployment of new infrastructure to achieve the energy transition.
@@ 2.1.2 Decoding of technical features and translation into MGA parameters
In practice, this leads to six (M=5, +1) different MGA runs that can be dealt with in parallel, similar to the original SPORES workflow. The number of designs to be generated within each of these parallel runs can be fine-tuned to match the same size of the original design space. In this study, we generate N=45 design options per parallel run, further distributing these 45 design options in three parallel batches of 15 designs each, in which we adopt different MGA weights ($w_{ij}$) to further increase diversity as done in previous work [lombardi_what_2023]. More precisely, we adopt, in one case, a simple integer weighting method; in another, a method based on deployment relative to maximum potential; and in the last case, a method based on the distance from the evolving average weight (see the Supplementary Methods for further details). In total, we obtain a design space of 270 system designs, slightly larger than the original one.
@@ 2.2 Model of the Portuguese energy system
The model accounts for Portugal’s entire energy demand across all sectors, spatially distributed between two macro-nodes, North and South. The model’s task is to satisfy the demand based on carbon-neutral energy technologies, which broadly consist of renewables, sustainable biofuels, and green hydrogen and derived e-fuels. The two nodes are connected via electricity transmission based on the network topology available from the Sector-Coupled Euro-Calliope. Within each node, however, renewable power capacity can be deployed in several sub-regions (18 in total), corresponding to the country’s administrative regions, with different land availability and capacity factors. Countries outside of Portugal are eliminated from the model and substituted with a simple representation of import and export options for electricity and hydrogen at a fixed price. For electricity, the price is set based on the historical average electricity price for trade with neighbouring Spain. For hydrogen, the price per kWh is calculated assuming a market price of 1.5 EUR/kg for green hydrogen in 2050 and 33.3 kWh/kg as the lower heating value. In this study, we use 2016 as the reference weather year, since we quantified it as a ‘typical’ year for Portugal within those available from the Euro-Calliope dataset.
@@ 3.1 The human-trained designs better align with high-level preferences in most cases
As anticipated in section 1, the new design space incorporates human preferences and knowledge of the problem, so we refer to it as the ‘human-trained design space’. It results from feeding the SPORES MGA algorithm with the decoded technical features that represent a high-level preference, plus all the (non-mutually-exclusive) features decoded across all preferences simultaneously. This produces a richer set of designs that align with the given high-level preferences, as can be seen in the comparison shown in Figure 4. In particular, the human-trained design space has substantially more designs matching those high-level preferences with a more direct connection to the deployment of specific technologies, such as the low use of hydrogen or the limited reliance on large-scale infrastructure that requires central planning. More nuanced preferences, such as the low regional concentration of wind farms, do improve, but less markedly. Finally, the high-level preference for a system design entailing a slow transition rate, i.e., a low overall deployment of new capacity, represents an exception for which the human-trained design space falls short in increasing the number of design options. This is a consequence of the method’s design, which relies on the intensification of features (Equation 3) to guide the search. These features are the capacities of technologies to be deployed. Thus, pushing technology capacities to their low or high extremes, especially when not diluted by doing so for multiple features at once, easily leads to more solutions with an overall high deployment of capacity, compared to the original design space that also entailed 80 solutions generated without any intensification (see sub-section 2.1.1).
One option to increase the number of features decoded as distinctive and hence intensified (i.e. minimised or maximised) simultaneously, is to reduce the distance-from-the-mean threshold for decoding. Figure S1 shows the outcome of the decoding procedure for a reduced (from 15% to 10%) threshold, which results in the design space labelled as ‘stricter’ in Figure 4. As expected, in this case, the percentage of designs with a slow transition rate improves compared to Figure 4, although it is still marginally less than in the original design space. Moreover, the percentage of designs aligned with a preference for limited central planning decreases, due to the ‘noise’ introduced by the additional features to be minimised or maximised alongside the push for less-centralised technologies, such as rooftop solar PV. There is hence a trade-off in having a stricter or more relaxed threshold for the decoding of distinctive technical features.
@@ 3.2 The potential for consensus formation increases substantially
Figure 6: Spatial deployment of renewable energy generation capacity, system-wide deployment of electrolysers and degree of reliance on imports in the system’s primary energy supply. The results are shown for two illustrative system design options, namely the options – in the original and in the human-trained (HITL-MGA) option spaces – with the best performance in terms of limited regional concentration of wind farms among the subset of overall best-performing, or ‘near-highest-consensus’, options.
@@ Conclusion
With this work, we set out to investigate whether an innovative human-in-the-loop approach to MGA could generate energy system design options that better align with stakeholder preferences and facilitate consensus formation compared to conventional, modeller-driven MGA. The results obtained by our original HITL-MGA workflow for a case study of the Portuguese energy system under controlled conditions (i.e., synthetic stakeholder preferences) show that this is the case, while also highlighting areas for further development.
@@
Moreover, we show that our HITL-MGA workflow effectively generates more designs that match the initial hidden high-level stakeholder preferences. Nonetheless, as an exception, we also detect that the intensification of features envisioned by our HITL-MGA workflow is structurally bound to conflict with potential high-level preferences for limited overall infrastructure deployment. In other words, intensifying a given technical feature comes with a substantial deployment of the related infrastructure. In our experiment for the case of Portugal, the only high-level preference for which the human-trained design space does not produce an improvement is the preference for a low rate of infrastructure deployment, which in our model setup coincides with less deployment overall. This is not surprising, since the nature of the energy system model used here is to model the deployment of technologies, and we use intensification (maximisation or minimisation) of technology capacity as the way to decode stakeholder preferences. Future work should examine the use of simple additional statistical measures to detect a preference for limited overall deployment and ways to translate that into the guided search. A possible translation could be reducing the strength of the b and c coefficients in Equation 3, representing the intensification of desired and undesired features, relative to coefficient a, aiming at the diversification of the system design. In previous work, we have already shown that such a manipulation effectively mitigates the degree of intensification of technical features [lombardi_what_2023].
@@
We also observe that the need to define a threshold for the automated decoding of statistically distinctive technical features in a given top-rated design entails a certain degree of uncertainty in the outcome of the decoding and, in turn, the human-trained design space. A tighter threshold, which we also examined, results in the decoding step detecting more features as distinctive, spreading and diluting the intensification objectives in Equation 3 across more features. This may work in favour of some high-level preferences, such as the one for limited infrastructure deployment overall, but less so for others. In general, the exact value of the threshold cannot be easily connected to any objective measure and, therefore, should be tested and calibrated beforehand, for instance, by means of a controlled experiment like the one we carry out in this study, in such a way as to match the needs of the model and research questions at stake.
@@
Regarding the likelihood of consensus formation, which we approximate via a simple MCDA analysis aligned with the assumed stakeholder preferences, the HITL-MGA workflow largely outperforms conventional MGA. In particular, the share of designs with an MCDA score within 25% of the highest score achieved in either design space increases from 1% to 18% when human inputs are integrated into the MGA loop. This is because the major improvements recorded in the number of stakeholder-appealing designs across most high-level preferences substantially counterpoise the marginally lower number of designs matching the ‘low-rate of deployment’ preference discussed above. In practice, the human-trained design space provides design options that allow stakeholders to compromise between their conflicting preferences, which were entirely missed by the original design space. For instance, designs that significantly limit the regional concentration and the overall deployment of onshore wind power capacity while simultaneously performing well on all the other assumed high-level preferences.
It is also worth noting that while we run the HITL-MGA workflow to decode and implement back into the MGA search system-wide technical features only, high-level preferences may well refer to specific features at the local scale. For instance, some stakeholders may be against a certain technology in a certain region, such as the deployment of electrolysers in regions affected by water scarcity, despite not opposing that technology per se. Even though we have not looked into this possibility in this first controlled experiment, nothing prevents modellers from running the same workflow for spatially explicit variables. Future work may further explore this possibility and its potential advantages.
@@
Overall, this study is the first to propose and successfully demonstrate a human-in-the-loop approach to MGA. Future work should explore the possible improvements of the method outlined above and test it further in a ‘field’ experiment with real-world stakeholders. The results of our controlled experiment show that HITL-MGA is a promising way to ensure that the generation of system design options matches stakeholder needs and facilitates consensus formation, thereby enabling and accelerating a just practical implementation of the energy transition.

### v2

@@
Abstract: The common use of cost minimisation to support energy system design decisions hides from view many economically comparable design options that stakeholders may prefer. Modelling to generate alternatives (MGA) is increasingly popular as a way to go beyond least-cost designs, providing stakeholders with diverse portfolios to appraise. However, generating all the feasible designs is not computationally viable; modellers must choose what design features to generate diversity around, despite not knowing which trade-offs matter the most in practice. Therefore, MGA alone cannot ensure the generation of design options that match stakeholder needs. To address this shortcoming, we propose a human-in-the-loop (HITL) approach that automatically integrates stakeholder preferences into MGA. We elicit preferences by letting stakeholders interact with a tentative MGA design space. Hence, we decode those preferences to feed them back to the MGA algorithm and perform a guided search. This search produces a human-trained design space with more designs that mirror the elicited preferences. A synthetic experiment for the Portuguese energy system shows that HITL-MGA may facilitate consensus formation, promising to accelerate technically and socially feasible energy transition decisions.
@@ 1 Introduction
Despite the urgency of climate mitigation, progress on the energy transition remains insufficient [iea_breakthrough_2023]. One of the key reasons for such insufficient progress is the difficulty in finding carbon-neutral energy system designs that are both technically feasible and acceptable for the many involved stakeholders, from citizens to industry, system operators, and policymakers [susser_why_2022]. Energy system optimisation models are increasingly looked at as a way to support these complex planning decisions, but severe limitations hamper their effectiveness. In particular, previous work has shown that the conventional practice of optimising the system design for cost hides from view many other feasible designs that are economically comparable – i.e., they are marginally more costly, or within the cost uncertainty range – while potentially having other un-modelled benefits that make them substantially more practically viable [lombardi_policy_2020].
@@
Approaches known as ‘Modelling to Generate Alternatives’, or MGA, first developed in the late 70s [brill_use_1979] make it possible to go beyond few, cost-optimal solutions and enlarge the decision space to hundreds of different possibilities, giving room for stakeholders to discuss and identify a consensus solution [decarolis_using_2011, decarolis_modelling_2016]. They have thus seen a spike in popularity: in the last few years, novel versions of MGA have been proposed for the specific needs of state-of-the-art large-scale energy system optimisation models that, representing large geographical regions at a high spatial resolution, entail countless degrees of spatial and technological freedom to achieve a feasible system design. Such novel MGA versions aim, for instance, at generating distinct spatial configurations of infrastructure deployment [lombardi_policy_2020] or at providing more homogeneous samples of the very large space of feasible decisions [pedersen_modeling_2021, neumann_broad_2023]. Increasingly, MGA is recognised as a critical tool to deliver meaningful, more transparent results [pickering_diversity_2022]; however, there are computational trade-offs in the use of MGA. For models of very large size, it is not yet computationally affordable to generate all the relevant design alternatives, and MGA algorithms need to decide what kind of diversity to prioritise in the design space. For instance, they must decide whether to focus more on discovering alternative technology mixes or different spatial configurations of technology deployment, for a given available amount of computational power [lombardi_what_2023]. And yet, determining which design options are more worth focussing on, insofar as they may be more relevant than others in facilitating real-world consensus formation, is not trivial and should not be dealt with by modellers alone to avoid biases [baker_just_2022].
[…]
@@
To test the above hypothesis, we perform an experiment under synthetic conditions in a case study of the Portuguese energy system. For our experiment, we rely on the SPORES MGA algorithm [lombardi_policy_2020], which lends itself to the proposed guided search due to its high degree of customisability, as we highlighted in previous work [lombardi_what_2023]. The synthetic conditions, represented by our use of synthetic stakeholder preferences, allow us to have a clear benchmark to quantitatively assess the relative merits of the updated, human-trained design space in terms of its degree of alignment with the assumed preferences and its capacity to facilitate consensus formation. Our results substantiate that the novel HITL-MGA approach comes with the expected advantages, and shows promise for generalisability and real-world application as a means to support the acceleration of technically and socially feasible energy planning decisions.
@@ 2 Methods
2.1 HITL-MGA workflow and synthetic experiment setup
@@
Figure 1: Conceptual representation of the proposed HITL-MGA workflow and of the synthetic experiment we carry out in this study to test the workflow. The theoretical framework we propose envisions that we collect the high-level system design preferences arising from a first iteration of an MGA design space with stakeholders. Then, we decode which technical features underlying such preferences and translate them into parameters that may inform and re-tune the MGA search for design options. The resulting updated design space is richer in design options that align with stakeholders’ high-level preferences, and it facilitates consensus formation. In our synthetic experiment, we replace stakeholder interaction with synthetic high-level stakeholder preferences, from which we derive five top-rated designs to initiate the HITL-MGA workflow. This way, we can test the outputs of the HITL-MGA workflow against the initial high-level preferences in a way that would not be possible if interacting with real-world stakeholders whose high-level preferences are typically unknown.
@@
In this study’s synthetic experiment, we test the method based on synthetic high-level stakeholder preferences, which allows us to benchmark the outputs in a way that would not be possible if interacting with real-world stakeholders whose high-level preferences, as anticipated above, may be unknown. We assume five high-level preferences:
@@
    A limited overall deployment of new infrastructure to achieve the energy transition.
@@ 2.1.2 Decoding of technical features and translation into MGA parameters
In practice, this leads to six (M=5, +1) different MGA runs that can be dealt with in parallel, similar to the original SPORES workflow. The number of designs to be generated within each of these parallel runs can be fine-tuned to match the same size of the original design space. In this study, we generate N=45 design options per parallel run, further distributing these 45 design options in three parallel batches of 15 designs each, in which we adopt different MGA weights ($w_{ij}$) to further increase diversity as done in previous work [lombardi_what_2023]. More precisely, we adopt, in one case, a simple integer weighting method; in another, a method based on deployment relative to maximum potential; and in the last case, a method based on the distance from the evolving average weight (see the Supporting Methods for further details). In total, we obtain a design space of 270 system designs, slightly larger than the original one. This similarity in size between the original and the HITL-MGA option spaces is purposeful so that we can assess potential benefits for a comparable computational effort. Both the original and the HITL-MGA option space are computed on the DelftBlue high-performance computer [DHPC2024].
@@ 2.2 Model of the Portuguese energy system
The model accounts for Portugal’s entire energy demand across all sectors, spatially distributed between two macro-nodes, North and South. The model’s task is to design the system in such a way as to satisfy the demand for a full year of operation based on carbon-neutral energy technologies, which broadly consist of renewables, sustainable biofuels, and green hydrogen and derived e-fuels, whose costs take into account projections towards 2050. The model only considers the final system design, or a ‘snapshot’ of what the system could look like in 2050, disregarding capacity deployment pathways through time on the way to such a final design. The two macro-nodes of North and South are connected via electricity transmission based on the network topology available from the Sector-Coupled Euro-Calliope. Within each node, however, renewable power capacity can be deployed in several sub-regions (18 in total), corresponding to the country’s administrative regions, with different land availability and capacity factors. Countries outside of Portugal are eliminated from the model and substituted with a simple representation of import and export options for electricity and hydrogen at a fixed price. For electricity, the price is set based on the historical average electricity price for trade with neighbouring Spain. For hydrogen, the price per kWh is calculated assuming a market price of 1.5 EUR/kg for green hydrogen in 2050 and 33.3 kWh/kg as the lower heating value. In this study, we use 2016 as the reference weather year, since we quantified it as a ‘typical’ year for Portugal within those available from the Euro-Calliope dataset.
@@ 3.1 The human-trained designs better align with high-level preferences in most cases
As anticipated in section 1, the new design space incorporates human preferences and knowledge of the problem, so we refer to it as the ‘human-trained design space’. It results from feeding the SPORES MGA algorithm with the decoded technical features that represent a high-level preference, plus all the (non-mutually-exclusive) features decoded across all preferences simultaneously. This produces a richer set of designs that align with the given high-level preferences, as can be seen in the comparison shown in Figure 4. In particular, the human-trained design space has substantially more designs matching those high-level preferences with a more direct connection to the deployment of specific technologies, such as the low use of hydrogen or the limited reliance on large-scale infrastructure that requires central planning. More nuanced preferences, such as the low regional concentration of wind farms, do improve, but less markedly. Finally, the high-level preference for a system design entailing a limited overall deployment of new infrastructure, represents an exception for which the human-trained design space falls short in increasing the number of design options. This is a consequence of the method’s design, which relies on the intensification of features (Equation 3) to guide the search. These features are the capacities of technologies to be deployed. Thus, pushing technology capacities to their low or high extremes, especially when not diluted by doing so for multiple features at once, easily leads to more solutions with an overall high deployment of capacity, compared to the original design space that also entailed 80 solutions generated without any intensification (see sub-section 2.1.1).
One option to increase the number of features decoded as distinctive and hence intensified (i.e. minimised or maximised) simultaneously, is to reduce the distance-from-the-mean threshold for decoding. Figure S1 shows the outcome of the decoding procedure for a reduced (from 15% to 10%) threshold, which results in the design space labelled as ‘stricter’ in Figure 4. As expected, in this case, the percentage of designs with limited overall deployment of new infrastructure improves compared to Figure 4, although it is still marginally less than in the original design space. Moreover, the percentage of designs aligned with a preference for limited central planning decreases, due to the ‘noise’ introduced by the additional features to be minimised or maximised alongside the push for less-centralised technologies, such as rooftop solar PV. There is hence a trade-off in having a stricter or more relaxed threshold for the decoding of distinctive technical features.
@@ 3.2 The potential for consensus formation increases substantially
Figure 6: Spatial deployment of renewable energy generation capacity, system-wide deployment of electrolysers and degree of reliance on imports in the system’s primary energy supply. The results are shown for two illustrative system design options, namely the options – in the original and in the human-trained (HITL-MGA) option spaces – with the best performance in terms of limited regional concentration of wind farms among the subset of overall best-performing, or ‘near-highest-consensus’, options. The geodata source to produce the maps is GISCO-Eurostat [eurostat_nuts].
@@ Conclusion
With this work, we set out to investigate whether an innovative human-in-the-loop approach to MGA could generate energy system design options that better align with stakeholder preferences and facilitate consensus formation compared to conventional, modeller-driven MGA. The results obtained by our original HITL-MGA workflow for a case study of the Portuguese energy system under synthetic conditions (i.e., synthetic stakeholder preferences) show that this is the case, while also highlighting areas for further development.
@@
Moreover, we show that our HITL-MGA workflow effectively generates more designs that match the initial hidden high-level stakeholder preferences. Nonetheless, as an exception, we also detect that the intensification of features envisioned by our HITL-MGA workflow is structurally bound to conflict with potential high-level preferences for limited new infrastructure deployment. In other words, intensifying a given technical feature comes with a substantial deployment of the related infrastructure. In our experiment for the case of Portugal, the only high-level preference for which the human-trained design space does not produce an improvement is the preference for a low rate of infrastructure deployment, which in our model setup coincides with less deployment overall. This is not surprising, since the nature of the energy system model used here is to model the deployment of technologies, and we use intensification (maximisation or minimisation) of technology capacity as the way to decode stakeholder preferences. Future work should examine the use of simple additional statistical measures to detect a preference for limited new infrastructure and ways to translate that into the guided search. A possible translation could be reducing the strength of the b and c coefficients in Equation 3, representing the intensification of desired and undesired features, relative to coefficient a, aiming at the diversification of the system design. In previous work, we have already shown that such a manipulation effectively mitigates the degree of intensification of technical features [lombardi_what_2023].
@@
We also observe that the need to define a threshold for the automated decoding of statistically distinctive technical features in a given top-rated design entails a certain degree of uncertainty in the outcome of the decoding and, in turn, the human-trained design space. A tighter threshold, which we also examined, results in the decoding step detecting more features as distinctive, spreading and diluting the intensification objectives in Equation 3 across more features. This may work in favour of some high-level preferences, such as the one for limited infrastructure deployment overall, but less so for others. In general, the exact value of the threshold cannot be easily connected to any objective measure and, therefore, should be tested and calibrated beforehand, for instance, by means of a synthetic experiment like the one we carry out in this study, in such a way as to match the needs of the model and research questions at stake.
@@
Regarding the likelihood of consensus formation, which we approximate via a simple MCDA analysis aligned with the assumed stakeholder preferences, the HITL-MGA workflow largely outperforms conventional MGA. In particular, the share of designs with an MCDA score within 25% of the highest score achieved in either design space increases from 1% to 18% when human inputs are integrated into the MGA loop. This is because the major improvements recorded in the number of stakeholder-appealing designs across most high-level preferences substantially counterpoise the marginally lower number of designs matching the ‘low-rate of deployment’ preference discussed above. In practice, the human-trained design space provides design options that allow stakeholders to compromise between their conflicting preferences, which were entirely missed by the original design space. For instance, designs that significantly limit the regional concentration and the overall deployment of onshore wind power capacity while simultaneously performing well on all the other assumed high-level preferences. Nonetheless, it is critical to underscore how, in a real-world setting, a computational approach that generates more middle-ground options by incorporating and synthesising multiple human inputs is an enabler but not a guarantee of the higher likelihood of consensus formation [royston_masters_2023]. It is paramount to complement the computational side with a collaboration with experts in stakeholder engagement and participatory action research that may ensure appropriate representation of all the stakeholder groups and a constructive deliberation process [mcgookin_participatory_2021, susser_rethink_2024].
In our synthetic experiment, we purposefully assume that the top-rated designs in the initial MGA space match some immediately understandable and not multifaceted preferences because this facilitates our testing and assessment of the improvements brought about by the HITL-MGA workflow. However, the workflow is entirely agnostic to the initial preferences and does not require any understanding of the rationale behind them. It only requires a list of the most-selected designs, which makes it applicable to any set of multifaceted real-world preferences. It is also worth noting that, while we run the HITL-MGA workflow to decode and implement back into the MGA search system-wide technical features only, high-level preferences may well refer to specific features at the local scale. For instance, some stakeholders may be against a certain technology in a certain region, such as the deployment of electrolysers in regions affected by water scarcity, despite not opposing that technology per se. Even though we have not looked into this possibility in this first synthetic experiment, nothing prevents modellers from running the same workflow for spatially explicit variables. Future work may further explore this possibility and its potential advantages. Another interesting possibility that future work may explore is the discretisation of the degree of (un)desirability of those distinctive technical features underlying a given top-rated design. Rather than simply flagging a feature as (un)desired, the method could accommodate incremental weights to further differentiate between highly or only moderately (un)desired features. Such discretisation would require a decision on how to define the discrete degrees of (un)desirability, further contributing to the need for calibration discussed earlier for the current threshold for feature distinctiveness. At the same time, it may further improve the accommodating of conflicting stakeholder preferences and the identification of middle-ground options. Finally, the HITL-MGA workflow in this study leverages the latest developments in terms of computationally efficient generation of MGA design spaces. However, the MGA field is continuously and rapidly evolving, and the proposed HITL-MGA would only benefit from any further computational advancements. Promising avenues for the further enhancement of HITL-MGA may be techniques such as the ex-post dimensionality reduction and metric interpolation proposed by a recent pre-print [lau_mgca_2024_preprint] or artificial intelligence techniques that are already common in the sister field of energy system multi-objective optimisation [prina_machine_2024].
The improved consensus-formation capabilities enabled by the HITL-MGA approach, and its very nature of integrating human preferences and knowledge of the problem into a computational workflow, align with the increasing demands for a tighter cross-fertilisation between engineering and other disciplines in tackling energy transition challenges at the interface between policy and science [baker_just_2022, mcgookin_advancing_2024]. Our human-trained search for design options answers the call for more ‘fair and unbiased algorithms’ [baker_just_2022] supporting energy transition decisions. And it does so by leveraging the principles of co-creation [bruckmann_towards_2023, mcgookin_advancing_2024] and providing a computational platform for their direct application.
@@
Overall, this study is the first to propose and successfully test a human-in-the-loop approach to MGA under synthetic conditions. Future work should explore the possible improvements of the method outlined above and test it further in a ‘field’ experiment with real-world stakeholders. The results of our synthetic experiment show that HITL-MGA is a promising way to ensure that the generation of system design options matches stakeholder needs and facilitates consensus formation, thereby enabling and accelerating a just practical implementation of the energy transition.